# Turbocharging Gaussian Process Inference with Approximate Sketch-and-Project

### v1

@@ B.2 Convergence along $-\ell$ dimensional subspace for positive-definite matrices
We start by introducing the notation and some results from existing literature for SAP and DPPs. Let $A$ be an $n\times n$ symmetric positive-definite matrix and let $A=VDV^{T}$ be its eigendecomposition, where $D$ is diagonal with entries $\lambda_{1}\geq...\geq\lambda_{n}$. For other matrices we use $\lambda_{i}(M)$ to denote the eigenvalues of $M$. For any vector $v\in\mathbb{R}^{n}$, define $\|v\|_{A}:=\sqrt{v^{T}Av}$. Let $V_{\ell}\in\mathbb{R}^{n\times\ell}$ consist of the $-\ell$ orthonormal eigenvectors of $A$ and $Q_{\ell}=V_{\ell}V_{\ell}^{T}$ be the orthogonal projector corresponding to the $-\ell$ eigenvectors of $A$. We rely on subsampling from fixed-size DPPs [Kulesza and Taskar, 2012, Dereziński and Mahoney, 2021] for our theoretical results. Here, for convenience, we consider the block size to be $2b$, in contrast to $b$ as considered in Algorithm 1. For a fixed $b$ and a positive-definite matrix $A$, we identify $2b$-DPP($A$) as a distribution over subsets of $\{1,\ldots,n\}$ of size $2b$, where any subset $\mathcal{B}$ has probability proportional to the corresponding principal submatrix of $A$, i.e., $\det(A_{\mathcal{B},\mathcal{B}})$. The subsampling matrix $S\in\mathbb{R}^{2b\times n}$ corresponding to the set $\mathcal{B}$ is defined as a matrix whose rows are the standard basis vectors associated with the indices in $\mathcal{B}$. For notational convenience, we will sometimes say that $S$ is drawn from $2b$-DPP($A$), or $S\sim 2b$-DPP($A$), keeping the index set $\mathcal{B}$ implicit.
@@
Lemma B.1 (Expected projection under $2b$-DPP($A$), adapted from Lemma 4.1 [Dereziński and Yang, 2024]).
Let $A\in\mathbb{R}^{n\times n}$ be a symmetric positive-definite matrix with eigenvalues $\lambda_{1}\geq\lambda_{2}\geq...\geq\lambda_{n}$, and let $21\leq b<n/$ 2 be fixed. Let $S\sim 2b$-DPP($A$). Then, we have
@@
Lemma B.2 (Sampling from $b$-DPP($A$), adapted from Anari et al. [2024]).
Given a positive-definite matrix $A$, there exists an algorithm that draws $t\geq 1$ approximate samples from $2b$-DPP($A$) in time $O(nb^{2}\log^{4}n+tb^{3}\log^{3}n)$. Furthermore, each drawn sample is indistinguishable from an exact sample from $2b$-DPP($A$) with probability at least $1-n^{-O(1)}$.
We now provide the main result of this section, incorporating the above two results in the analysis of sketch-and-project (SAP) along the $-\ell$ subspace of $A$. Since throughout this section we focus on a linear system with a single right-hand side, we will use $w_{t}$ instead of $W_{t}$ to denote the SAP iterates, as these are always vectors. The update rule for SAP is now:
@@
Theorem B.3 (Fast convergence along $-\ell$ subspace).
Let $b<n/2$ be fixed and $S\in\mathbb{R}^{2b\times n}$ be a random subsampling matrix sampled from $2b$-DPP($A$). Furthermore, let $1\leq\ell<b$ be also fixed and $\lambda_{\ell}(\bar{\Pi})\geq 2\lambda_{n}(\bar{\Pi})$. For any $t>1$ define $\bar{w}_{t}=\frac{2}{t}\sum_{i=t/2}^{t-1}{w_{i}}$ where $w_{i}$ are generated using (5). Then,
@@
The assumption $\lambda_{\ell}(\bar{\Pi})\geq 2\lambda_{n}(\bar{\Pi})$ is not restrictive buts lets us provide a cleaner analysis by avoiding the edge case of $\lambda_{\ell}(\bar{\Pi})\approx\lambda_{n}(\bar{\Pi})$. In fact, in this corner case, we can simply rely on the existing SAP analysis from previous works and recover our main result, Theorem 3.2. For completeness, we derive Theorem B.9 for posterior GP mean inference in Section B.3 without the assumption $\lambda_{\ell}(\bar{\Pi})\geq 2\lambda_{n}(\bar{\Pi})$. The proof of Theorem B.3 appears after the proof of Theorem B.8. We build towards the proof starting with the following result for SAP [Dereziński and Yang, 2024]:
@@
Let $S\in\mathbb{R}^{2b\times n}$ be a random subsampling matrix sampled from $2b$-DPP($A$). Then,
@@
However, in the following result, we show that along the $-\ell$ eigenspace of $A$, the expected iterates can converge at a much faster rate than $1-\lambda_{n}(\bar{\Pi})$.
Lemma B.5 (Convergence of expected iterates along $-\ell$ subspace).
Let $S\in\mathbb{R}^{2b\times n}$ be a random subsampling matrix sampled from $2b$-DPP($A$). Then
@@
As $A=VDV^{T}$ we have $\Pi=VD^{1/2}V^{T}S^{T}(SVDV^{T}S^{T})^{\dagger}SVD^{1/2}V^{T}$. Due to Lemma B.1 we know that $\bar{\Pi}$ and $A$ share the same eigenvectors, therefore, $\bar{\Pi}$ and $Q_{\ell}$ commute. Consequently,
@@
We now analyze convergence along the $-\ell$ dimensional subspace in L2-norm by considering $\mathbb{E}\|A^{1/2}(w_{t+1}-w_{\star})\|^{2}_{Q_{\ell}}$. We have,
@@
Lemma B.6 (L2-norm error along $-\ell$ subspace).
Let $S\in\mathbb{R}^{2b\times n}$ be a random subsampling matrix sampled from $2b$-DPP($A$). Then
@@
Crucially, the first two terms terms in (6) live in the $-\ell$ subspace, but the third term does not. Nevertheless, we are still able to bound it as follows:
@@
Define $\gamma=1-\lambda_{n}(\bar{\Pi})$. Invoking Lemma B.6 we deduce,
@@
Now, recursively applying the tower rule and Lemma B.4 yields
@@
We now use the tail averaging idea similar to Epperly et al. [2024a], obtaining fast convergence along the $-\ell$ subspace. The proof of Theorem B.3 is then derived from the following result, after combining it with fixed-size DPP sampling guarantees of Lemma B.1.
@@
Let $S\in\mathbb{R}^{2b\times n}$ be a random subsampling matrix sampled from $2b$-DPP($A$). For any $t>2$ define $\hat{w}_{t}=\frac{2}{t}\sum_{i\geq t/2}^{t-1}{w_{i}}$ where $w_{i}$ are generated using update rule (5). Then
@@
Applying Lemma B.6 to upper bound $\mathbb{E}\|A^{1/2}(w_{r}-w_{\star})\|^{2}_{Q_{\ell}}$ obtains
@@
Completing the proof of Theorem B.3
@@
Recalling the definition of $\alpha$ from Theorem B.8 and our assumption that $\lambda_{\ell}(\bar{\Pi})\geq 2\lambda_{n}(\bar{\Pi})$, we obtain for $\gamma=1-\lambda_{n}(\bar{\Pi})$ that $(1-\alpha)^{-1}<2\gamma/\lambda_{\ell}(\bar{\Pi})$. Consequently,
@@
<$\overset{(1)}{<}4\left(\frac{1}{b}\sum_{i>b}\frac{\lambda_{i}}{\lambda_{\ell}}\right)\gamma$
=$\overset{(2)}{=}4\phi(b,\ell)\gamma,$
where $(1)$ applies Lemma B.1 and $(2)$ defines $\phi(b,\ell)=\frac{1}{b}\sum_{i>b}\frac{\lambda_{i}}{\lambda_{\ell}}$. Observing the elementary inequalities:
@@
we immediately deduce from Theorem B.8 that
@@
B.3 Posterior mean inference along $-\ell$ subspace for GPs
@@
In particular, $m_{n}=\sum_{i=1}^{n}{(w_{\star})_{i}k(\cdot,x_{i})}$, where $w_{\star}=(K+\lambda I)^{-1}y$. Furthermore, note that for any element $m^{\prime}\in\mathcal{H}_{n}$, we have $\|m^{\prime}\|_{\mathcal{H}}^{2}=w^{T}Kw=\|w\|_{K}^{2}$. The operator $\mathcal{C}_{n}:=\frac{1}{n}\sum_{i=1}^{n}{k(\cdot,x_{i})\otimes k(\cdot,x_{i})}$ is known as the empirical covariance operator. Let $v_{j}$ denote the $j^{th}$ unit eigenvector of $K$ with eigenvalue $\lambda_{j}$. It is straightforward to show that $u_{j}:=\frac{1}{\sqrt{\lambda_{j}}}\sum_{i=1}^{n}{v_{ji}k(\cdot,x_{i})}$ is a unit eigenvector of the unnormalized empirical covariance operator $\sum_{i=1}^{n}{k(\cdot,x_{i})\otimes k(\cdot,x_{i})}$ with eigenvalue $\lambda_{j}$. Let $V_{\ell}\in\mathbb{R}^{n\times\ell}$ consists of $-\ell$ orthogonal eigenvectors of $K$ as columns and $Q_{\ell}=V_{\ell}V_{\ell}^{T}$ be a projection matrix onto the subspace spanned by $v_{1},\ldots,v_{\ell}$. Consider the $\ell$-dimensional subspace $\mathcal{H}_{\ell}$ defined as
:=$\mathcal{H}_{\ell}:=\left\{\sum_{i=1}^{n}{(Q_{\ell}w)_{i}k(\cdot,x_{i})}\ |\ w\in\mathbb{R}^{n}\right\}.$
We claim that $\mathcal{H}_{\ell}$ is the subspace formed by $-\ell$ eigenvectors of the empirical covariance operator. This can be seen clearly by choosing $w=v_{j}$ for $1\leq j\leq\ell$, we get $u_{j}\in\mathcal{H}_{\ell}$. We have the following conclusions.
@@
Let $b<n/2$ and $S\in\mathbb{R}^{2b\times n}$ be a random subsampling matrix sampled from $2b$-DPP($K_{\lambda}$). For any $t>2$ define $\hat{w}_{t}=\frac{2}{t}\sum_{i=t/2}^{t-1}{w_{i}}$ where $w_{i}$ are generated using the update rule (5). Then, sketch-and-project initialized at 00 satisfies
@@
as we assumed $b>\sum_{i>b}\frac{\lambda_{i}}{\lambda_{\ell}+\lambda}$. Furthermore, for given spectral decay for any $\beta>1$, we have $\ell+\sum_{i>b}\frac{\lambda_{i}}{\lambda_{\ell}}<2\ell$. This implies for $b=2\ell$ we obtain $\epsilon$ accuracy result in $t=\mathcal{O}\left(\frac{n}{\ell}\min\left\{\frac{\log(n/\ell)}{\epsilon},\left(1+\frac{\ell}{n}\frac{\lambda_{\ell}+\lambda}{\lambda_{n}+\lambda}\right)\log(n/\ell\epsilon)\right\}\right)$, where we used $\lambda=\mathcal{O}(1)$ to get $\log\left(\frac{n}{\ell}(1+\lambda/\lambda_{\ell})\right)=\mathcal{O}\left(\log(\frac{n}{\ell})\right)$. Then, after applying Markov’s inequality, $\hat{m}=\hat{m}_{t}$ obtains the desired $\epsilon$ accuracy with probability 0.999. The total time complexity follows by combining the number of iterations with cost of approximate sampling from $2b$-DPP($A$) (Lemma B.2) and additional per iteration cost of $\mathcal{O}(n\ell+\ell^{3}\log^{3}n)$. Note that while the sampling algorithm is not exact, taking the union bound with respect to the high probability guarantees in Lemma B.2 ensures that all of the samples are indistinguishable from true DPP with probability 0.999. Finally, union bounding over the three 0.999 probability events we have invoked concludes the proof. ∎
@@
We reconsider case 1 from the proof of Corollary B.10. Using the given spectral decay profile for $K$ and that $\lambda<\frac{1}{nb^{\beta-1}}$, we get $\phi(b,\ell)=\mathcal{O}\left(\frac{b^{-\beta}}{\ell^{-\beta}}\right)$. Therefore after $t=\frac{n}{\ell\epsilon}\frac{b^{-\beta}}{\ell^{-\beta}}$ we obtain the $\epsilon$ approximation guarantee. The total runtime complexity becomes:

### v2

@@ B.2 Convergence along top-$\ell$ dimensional subspace for positive-definite matrices
We start by introducing the notation and some results from existing literature for SAP and DPPs. Let $A$ be an $n\times n$ symmetric positive-definite matrix and let $A=VDV^{T}$ be its eigendecomposition, where $D$ is diagonal with entries $\lambda_{1}\geq...\geq\lambda_{n}$. For other matrices we use $\lambda_{i}(M)$ to denote the eigenvalues of $M$. For any vector $v\in\mathbb{R}^{n}$, define $\|v\|_{A}:=\sqrt{v^{T}Av}$. Let $V_{\ell}\in\mathbb{R}^{n\times\ell}$ consist of the top-$\ell$ orthonormal eigenvectors of $A$ and $Q_{\ell}=V_{\ell}V_{\ell}^{T}$ be the orthogonal projector corresponding to the top-$\ell$ eigenvectors of $A$. We rely on subsampling from fixed-size DPPs [Kulesza and Taskar, 2012, Dereziński and Mahoney, 2021] for our theoretical results. Here, for convenience, we consider the block size to be $2b$, in contrast to $b$ as considered in Algorithm˜1. For a fixed $b$ and a positive-definite matrix $A$, we identify $2b$-DPP(AA) as a distribution over subsets of $\{1,\ldots,n\}$ of size $2b$, where any subset $\mathcal{B}$ has probability proportional to the corresponding principal submatrix of $A$, i.e., $\det(A_{\mathcal{B},\mathcal{B}})$. The subsampling matrix $S\in\mathbb{R}^{2b\times n}$ corresponding to the set $\mathcal{B}$ is defined as a matrix whose rows are the standard basis vectors associated with the indices in $\mathcal{B}$. For notational convenience, we will sometimes say that $S$ is drawn from $2b$-DPP(AA), or $S\sim 2b$-DPP(AA), keeping the index set $\mathcal{B}$ implicit.
@@
Lemma B.1 (Expected projection under $2b$-DPP(AA), adapted from Lemma 4.1 [Dereziński and Yang, 2024]).
Let $A\in\mathbb{R}^{n\times n}$ be a symmetric positive-definite matrix with eigenvalues $\lambda_{1}\geq\lambda_{2}\geq...\geq\lambda_{n}$, and let $1\leq b<n/2$ be fixed. Let $S\sim 2b$-DPP(AA). Then, we have
@@
Lemma B.2 (Sampling from $b$-DPP(AA), adapted from Anari et al. [2024]).
Given a positive-definite matrix $A$, there exists an algorithm that draws $t\geq 1$ approximate samples from $2b$-DPP(AA) in time $O(nb^{2}\log^{4}n+tb^{3}\log^{3}n)$. Furthermore, each drawn sample is indistinguishable from an exact sample from $2b$-DPP(AA) with probability at least $1-n^{-O(1)}$.
We now provide the main result of this section, incorporating the above two results in the analysis of sketch-and-project (SAP) along the top-$\ell$ subspace of $A$. Since throughout this section we focus on a linear system with a single right-hand side, we will use $w_{t}$ instead of $W_{t}$ to denote the SAP iterates, as these are always vectors. The update rule for SAP is now:
@@
Theorem B.3 (Fast convergence along top-$\ell$ subspace).
Let $b<n/2$ be fixed and $S\in\mathbb{R}^{2b\times n}$ be a random subsampling matrix sampled from $2b$-DPP(AA). Furthermore, let $1\leq\ell<b$ be also fixed and $\lambda_{\ell}(\bar{\Pi})\geq 2\lambda_{n}(\bar{\Pi})$. For any $t>1$ define $\bar{w}_{t}=\frac{2}{t}\sum_{i=t/2}^{t-1}{w_{i}}$ where $w_{i}$ are generated using (5). Then,
@@
The assumption $\lambda_{\ell}(\bar{\Pi})\geq 2\lambda_{n}(\bar{\Pi})$ is not restrictive buts lets us provide a cleaner analysis by avoiding the edge case of $\lambda_{\ell}(\bar{\Pi})\approx\lambda_{n}(\bar{\Pi})$. In fact, in this corner case, we can simply rely on the existing SAP analysis from previous works and recover our main result, Theorem 3.2. For completeness, we derive Theorem B.9 for posterior GP mean inference in Section B.3 without the assumption $\lambda_{\ell}(\bar{\Pi})\geq 2\lambda_{n}(\bar{\Pi})$. The proof of Theorem˜B.3 appears after the proof of Theorem˜B.8. We build towards the proof starting with the following result for SAP [Dereziński and Yang, 2024]:
@@
Let $S\in\mathbb{R}^{2b\times n}$ be a random subsampling matrix sampled from $2b$-DPP(AA). Then,
@@
However, in the following result, we show that along the top-$\ell$ eigenspace of $A$, the expected iterates can converge at a much faster rate than $1-\lambda_{n}(\bar{\Pi})$.
Lemma B.5 (Convergence of expected iterates along top-$\ell$ subspace).
Let $S\in\mathbb{R}^{2b\times n}$ be a random subsampling matrix sampled from $2b$-DPP(AA). Then
@@
As $A=VDV^{T}$ we have $\Pi=VD^{1/2}V^{T}S^{T}(SVDV^{T}S^{T})^{\dagger}SVD^{1/2}V^{T}$. Due to Lemma˜B.1 we know that $\bar{\Pi}$ and $A$ share the same eigenvectors, therefore, $\bar{\Pi}$ and $Q_{\ell}$ commute. Consequently,
@@
We now analyze convergence along the top-$\ell$ dimensional subspace in L2-norm by considering $\mathbb{E}\|A^{1/2}(w_{t+1}-w_{\star})\|^{2}_{Q_{\ell}}$. We have,
@@
Lemma B.6 (L2-norm error along top-$\ell$ subspace).
Let $S\in\mathbb{R}^{2b\times n}$ be a random subsampling matrix sampled from $2b$-DPP(AA). Then
@@
Crucially, the first two terms terms in (6) live in the top-$\ell$ subspace, but the third term does not. Nevertheless, we are still able to bound it as follows:
@@
Define $\gamma=1-\lambda_{n}(\bar{\Pi})$. Invoking Lemma˜B.6 we deduce,
@@
Now, recursively applying the tower rule and Lemma˜B.4 yields
@@
We now use the tail averaging idea similar to Epperly et al. [2024a], obtaining fast convergence along the top-$\ell$ subspace. The proof of Theorem˜B.3 is then derived from the following result, after combining it with fixed-size DPP sampling guarantees of Lemma˜B.1.
@@
Let $S\in\mathbb{R}^{2b\times n}$ be a random subsampling matrix sampled from $2b$-DPP(AA). For any $t>2$ define $\hat{w}_{t}=\frac{2}{t}\sum_{i\geq t/2}^{t-1}{w_{i}}$ where $w_{i}$ are generated using update rule (5). Then
@@
Applying Lemma˜B.6 to upper bound $\mathbb{E}\|A^{1/2}(w_{r}-w_{\star})\|^{2}_{Q_{\ell}}$ obtains
@@
Completing the proof of Theorem˜B.3
@@
Recalling the definition of $\alpha$ from Theorem˜B.8 and our assumption that $\lambda_{\ell}(\bar{\Pi})\geq 2\lambda_{n}(\bar{\Pi})$, we obtain for $\gamma=1-\lambda_{n}(\bar{\Pi})$ that $(1-\alpha)^{-1}<2\gamma/\lambda_{\ell}(\bar{\Pi})$. Consequently,
@@
$\overset{(1)}{<}4\left(\frac{1}{b}\sum_{i>b}\frac{\lambda_{i}}{\lambda_{\ell}}\right)\gamma$
$\overset{(2)}{=}4\phi(b,\ell)\gamma,$
where $(1)$ applies Lemma˜B.1 and $(2)$ defines $\phi(b,\ell)=\frac{1}{b}\sum_{i>b}\frac{\lambda_{i}}{\lambda_{\ell}}$. Observing the elementary inequalities:
@@
we immediately deduce from Theorem˜B.8 that
@@
B.3 Posterior mean inference along top-$\ell$ subspace for GPs
@@
In particular, $m_{n}=\sum_{i=1}^{n}{(w_{\star})_{i}k(\cdot,x_{i})}$, where $w_{\star}=(K+\lambda I)^{-1}y$. Furthermore, note that for any element $m^{\prime}\in\mathcal{H}_{n}$, we have $\|m^{\prime}\|_{\mathcal{H}}^{2}=w^{T}Kw=\|w\|_{K}^{2}$. The operator $\mathcal{C}_{n}:=\frac{1}{n}\sum_{i=1}^{n}{k(\cdot,x_{i})\otimes k(\cdot,x_{i})}$ is known as the empirical covariance operator. Let $v_{j}$ denote the $j^{th}$ unit eigenvector of $K$ with eigenvalue $\lambda_{j}$. It is straightforward to show that $u_{j}:=\frac{1}{\sqrt{\lambda_{j}}}\sum_{i=1}^{n}{v_{ji}k(\cdot,x_{i})}$ is a unit eigenvector of the unnormalized empirical covariance operator $\sum_{i=1}^{n}{k(\cdot,x_{i})\otimes k(\cdot,x_{i})}$ with eigenvalue $\lambda_{j}$. Let $V_{\ell}\in\mathbb{R}^{n\times\ell}$ consists of top-$\ell$ orthogonal eigenvectors of $K$ as columns and $Q_{\ell}=V_{\ell}V_{\ell}^{T}$ be a projection matrix onto the subspace spanned by $v_{1},\ldots,v_{\ell}$. Consider the $\ell$-dimensional subspace $\mathcal{H}_{\ell}$ defined as
$\mathcal{H}_{\ell}:=\left\{\sum_{i=1}^{n}{(Q_{\ell}w)_{i}k(\cdot,x_{i})}\ |\ w\in\mathbb{R}^{n}\right\}.$
We claim that $\mathcal{H}_{\ell}$ is the subspace formed by top-$\ell$ eigenvectors of the empirical covariance operator. This can be seen clearly by choosing $w=v_{j}$ for $1\leq j\leq\ell$, we get $u_{j}\in\mathcal{H}_{\ell}$. We have the following conclusions.
@@
Let $b<n/2$ and $S\in\mathbb{R}^{2b\times n}$ be a random subsampling matrix sampled from $2b$-DPP($K_{\lambda}$). For any $t>2$ define $\hat{w}_{t}=\frac{2}{t}\sum_{i=t/2}^{t-1}{w_{i}}$ where $w_{i}$ are generated using the update rule (5). Then, sketch-and-project initialized at 0 satisfies
@@
as we assumed $b>\sum_{i>b}\frac{\lambda_{i}}{\lambda_{\ell}+\lambda}$. Furthermore, for given spectral decay for any $\beta>1$, we have $\ell+\sum_{i>b}\frac{\lambda_{i}}{\lambda_{\ell}}<2\ell$. This implies for $b=2\ell$ we obtain $\epsilon$ accuracy result in $t=\mathcal{O}\left(\frac{n}{\ell}\min\left\{\frac{\log(n/\ell)}{\epsilon},\left(1+\frac{\ell}{n}\frac{\lambda_{\ell}+\lambda}{\lambda_{n}+\lambda}\right)\log(n/\ell\epsilon)\right\}\right)$, where we used $\lambda=\mathcal{O}(1)$ to get $\log\left(\frac{n}{\ell}(1+\lambda/\lambda_{\ell})\right)=\mathcal{O}\left(\log(\frac{n}{\ell})\right)$. Then, after applying Markov’s inequality, $\hat{m}=\hat{m}_{t}$ obtains the desired $\epsilon$ accuracy with probability 0.999. The total time complexity follows by combining the number of iterations with cost of approximate sampling from $2b$-DPP(AA) (Lemma˜B.2) and additional per iteration cost of $\mathcal{O}(n\ell+\ell^{3}\log^{3}n)$. Note that while the sampling algorithm is not exact, taking the union bound with respect to the high probability guarantees in Lemma˜B.2 ensures that all of the samples are indistinguishable from true DPP with probability 0.999. Finally, union bounding over the three 0.999 probability events we have invoked concludes the proof. ∎
@@
We reconsider case 1 from the proof of Corollary˜B.10. Using the given spectral decay profile for $K$ and that $\lambda<\frac{1}{nb^{\beta-1}}$, we get $\phi(b,\ell)=\mathcal{O}\left(\frac{b^{-\beta}}{\ell^{-\beta}}\right)$. Therefore after $t=\frac{n}{\ell\epsilon}\frac{b^{-\beta}}{\ell^{-\beta}}$ we obtain the $\epsilon$ approximation guarantee. The total runtime complexity becomes: